OpenStudy (anonymous):

Determine the nature of the solutions of the equations (how many solutions): x^2-18x+81=0

OpenStudy (anonymous):

\[(x-9)^2\]

OpenStudy (anonymous):

x = 9

OpenStudy (jamesj):

Note that x^2 - 18x + 81 = (x - 9)^2. So this is equal to zero if and olnly if x - 9 = 0 , i.e., x = 9. Hence there is just one solution.

OpenStudy (anonymous):

only one real root x=9